Find two positive numbers so that twice their sum equals their product and one number is 8 times the other number. Enter the smaller number first.
Answer by checkley77(12569)

2(x+y)=xy
x=8y
2(8y+y)=8y*y
2*9y=8y^2
8y^2-18y=0
2y(4y-9)=0
4y-9=0
4y=9
y=9/4
y=2.25 ans.
x=8*2.25
x=18 ans.
Proof:
2*(18+2.25)=18*2.25
2*20.25=40.5
40.5=40.5